Assume that x has a normal distribution with the specified mean and standard deviation. Find the indicated probability. (Round your answer to four decimal places.)

μ = 2.9; σ = 0.35

P(x ≥ 2) =
Answer #1

soluton Given : ا 2,9 د3 0 : 0 (2 ی دام - : ( ج )م وه ، م-ا = 2 - 22 م - - =-P ( 3 ) = - Pa< - 25 7/4) ک3 اد0.00 -= - C,9949